Sharp Company manufactures a product for which the following standards have been set:

Standard Quantity
or Hours
Standard Price
or Rate
Standard
Cost
  Direct materials 3 feet $ 5 per foot $ 15
  Direct labor ? hours ? per hour ?

     During March, the company purchased direct materials at a cost of $45,375, all of which were used in the production of 2,350 units of product. In addition, 4,800 hours of direct labor time were worked on the product during the month. The cost of this labor time was $50,400. The following variances have been computed for the month:

  Materials quantity variance $ 2,250 U  
  Labor spending variance $ 3,400 U  
  Labor efficiency variance $ 1,000 U  
Required:
1. For direct materials:
a.

Compute the actual cost per foot for materials for March. (Round your answer to 2 decimal places.)

     

Actual cost $6.00 per foot
b.

Compute the price variance and the spending variance. (Do not round intermediate calculations. Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance))


Price variance $7,350 U

Spending variance$8,850 U

2.

For direct labor: (Do not round intermediate calculations.)

a.

Compute the standard direct labor rate per hour. (Round your final answer to 2 decimal places.)

     

Standard direct labor rate per hour $2,400.00

  

b.

Compute the standard hours allowed for the month’s production

Standard hours 4,700 hours

     

c.

Compute the standard hours allowed per unit of product. (Round your answer to 1 decimal place.)

Standard hours 2.0 hours per unit
Answer #1

Solution 1a:

Material quantity variance = $2,250 U

(SQ - AQ)*SP = -$2,250

(2350*3 - AQ) * $5 = - $2,250

Actual quantity = 7500 Foot

Actual cost of purchase = $45,375

actual cost per foot = $45,375 / 7500 = $6.05 per foot

Solution 1b:

Material price variance = (SP - AP) * AQ = ($5 - $6.05) * 7500 = $7,875 U

Material spending variance = MPV + MUV = $7,875 U + $2,250 U = $10,125 U

Solution 2a:

Actual rate of labor per hour = $50,400 / 4800 = $10.50 per hour

Labor rate variance = Labor spending variance - Labor efficiency variance = $3,400 U - $1,000 U = $2,400 U

(SR - AR) * AH = - $2,400

(SR - $10.50) * 4800 = - $2,400

SR = $10 per hour

Solution 2b:

Labor efficiency variance = (SH - AH) * SR = (SH - 4800) * $10 = - $1,000

Standard hours = 4700 hours

solution 2c:

Standard hours allowed per unit = 4700 / 2350 = 2 hour per unit.